OpenStudy (anonymous):

If f(x)=x^3+4x+6, find (f−1)′(6)

OpenStudy (anonymous):

its actaully( f^-1)'(6)

OpenStudy (anonymous):

i meant to correct that. sorry

myininaya (myininaya):

\[(f^{-1})'(6)=\frac{1}{f'(f^{-1}(6))}=\frac{1}{f'(0)} \text{ since } f(0)=6 => f^{-1}(6)=0\]

OpenStudy (anonymous):

0 doesnt work

OpenStudy (anonymous):

Can i have some help?

myininaya (myininaya):

\[f'(x)=3x^2+4 =>f'(0)=3(0)^2+4=4 => (f^{-1})'(6)=\frac{1}{4}\]

OpenStudy (anonymous):

Oh, he meant inverse. I'm sorry. I'll delete my answer.

OpenStudy (anonymous):

1/4 works thanks myinaya

myininaya (myininaya):

I know it works lol

myininaya (myininaya):

Oh you thought I was saying the answer is 0? Do you understand the steps?